# Millimeter-wave Radio SLAM: End-to-End Processing Methods and Experimental Validation

### v1

@@ Abstract
In this article, we address the timely topic of cellular bistatic simultaneous localization and mapping (SLAM) with specific focus on complete processing solutions from raw I/Q samples to user equipment (UE) and landmark location information in millimeter-wave (mmWave) networks. Firstly, we propose a new multipath channel parameter estimation solution which operates directly with beam reference signal received power (BRSRP) measurements, alleviating the need to know the true antenna beampatterns or the underlying beamforming weights. Additionally, the method has built-in robustness against unavoidable antenna sidelobes. Secondly, we propose new snapshot SLAM algorithms that have increased robustness and identifiability compared to prior-art, in practical built environments with complex clutter and multi-bounce propagation scenarios. The performance of the proposed methods is assessed at the 60 GHz mmWave band, via both realistic ray-tracing evaluations as well as true experimental measurements, in an indoor environment. Wide set of offered results clearly demonstrate the improved performance, compared to the relevant prior-art, in terms of the channel parameter estimation as well as the end-to-end SLAM performance. Finally, the article provides the measured 60 GHz data openly available for the research community, facilitating results reproducibility as well as further algorithm development.
@@ I Introduction
W hile the primary purpose of mobile cellular networks is to provide efficient connectivity services, the ability to extract location information and situational awareness of the surrounding environment is also receiving increasing interest [1, 2, 3, 4]. The related notion of integrated sensing and communications (ISAC) refers to extending the situational awareness from ordinary user equipment (UE) positioning to the ability to sense also various passive objects in the environment, through cellular radio-based measurements and corresponding signal processing [5, 1]. Such knowledge of the UE locations and surrounding environment can be harnessed in numerous ways, for example in different extended reality (XR) use cases [6], vehicular applications [7], or industrial systems [8]. In general, the prospects for high-accuracy situational awareness are known to improve [9, 10] when the networks are expanding towards the millimeter-wave (mmWave) frequency bands. The current fifth generation New Radio (5G NR) specifications support already operating bands up to 71 GHz [11], while further extensions towards the sub-THz regime are expected in the 6G era [12].
Bistatic cellular simultaneous localization and mapping (SLAM) is one of the most prominent, yet technically challenging ISAC scenarios where the coordinates of both the UE as well as those of the environment scattering points – commonly referred to as the landmarks – are all jointly estimated, based on either uplink (UL) or downlink (DL) reference signals and known base-station (BS) locations [13, 14, 15]. This is illustrated conceptually in Fig. 1. Complete end-to-end solutions for SLAM comprise estimators for spatio-temporal channel parameters, often in the form of time of arrival (ToA), angle-of-arrival (AoA), and/or angle-of-departure (AoD) of the involved propagation paths, combined with the actual SLAM filters or snapshot estimators that process the channel parameter estimates into the corresponding estimates of the UE and landmark locations [16]. This is also the main scope of this article. We have specific focus on implementation-feasible yet high-performing end-to-end processing solutions that can operate with minimal knowledge of the involved antenna system beampatterns and facilitate mmWave SLAM in practical complex built environments, particularly indoors, while operating on downlink positioning reference signal (PRS) standardized already for the existing 5G NR networks. Additionally, we emphasize research reproducibility and provide measured mmWave I/Q and channel data openly for the research community, while use the same measured data to evaluate and benchmark the proposed methods against the relevant prior art.
@@ I-A Prior Art
The available literature on channel parameter estimation is generally-speaking wide, however, a vast majority of works is carried out under the assumptions of known antenna steering vectors, beamforming weights, and thereon beampatterns. Such methods are described, e.g., in [17, 18, 19, 20, 21, 22, 23, 24, 25]. To this end, [17] and [18] present Bayesian channel estimation algorithms based on different variants of Kalman filters, while [19, 21, 22] harness the mmWave channel sparsity through compressed sensing (CS) methods. The work in [20], in turn, proposes an AoA estimation method using virtual subarrays which, however, requires a special beamformer or antenna pattern design, similar to [23]. Furthermore, [24, 25] propose joint AoD and AoA estimation methods with specific frequency-dependent codebooks in true-time-delay array context. Opposed to these previous works, an angle estimation algorithm building on the beam reference signal received power (BRSRP) measurements is described in the recent work in [15]. The method builds on thresholding and successive cancellation principle, operating on AoA-AoD power map, however, is lacking, e.g., explicit treatment of antenna sidelobes whose impact can be substantial in mmWave systems. This is particularly so in SLAM context where the multipaths and their dynamic range is one key factor. For clarity, we also note that works like [26] exist that deploy machine-learning to the channel parameters estimation – however, such works are still commonly at their early phases.
@@
In the context of the available SLAM methods, we first note that both snapshot approaches [16, 27, 28, 29] as well as sequential filtering based solutions [30, 13, 14, 15, 31] exist. Both research directions are relevant and the preferred choice depends on the overall system and application scenario. Snapshot SLAM is fundamentally important as it serves as a baseline for what can be done with radio signals alone, without any movement models, while a snapshot method can also be used as input to filtering [27]. On the other hand, filtering based SLAM methods process the observations sequentially over time and are expected to remove false detections and improve the accuracy [14]. Both SLAM methods also have their limitations. A major drawback of snapshot SLAM is that it is not always applicable, since the measurements may not be sufficient to solve the SLAM problem [16]. For example, the UE cannot be localized with the line-of-sight (LoS) alone if the clocks of the UE and BS are not synchronized [27] and one must resort to filtering based techniques to solve the problem over time. On the other hand, filtering methods always require a snapshot algorithm for initialization when prior information is not available. Another disadvantage of filtering methods is that they require solving a complex data association problem which increases computational overhead of the algorithms [14, 31]. It is to be noted that low complexity alternatives exist [30, 13], however, at the expense of reduced accuracy.
@@ I-B Novelty and Contributions
Compared to the available methods and literature, this article provides the following contributions. Firstly, opposed to the vast majority of existing literature in [17, 18, 19, 20, 21, 22, 23, 24, 25], we focus on enabling accurate multipath AoA/AoD estimation with BRSRP measurements only, without knowledge of the complex antenna patterns or the underlying steering vectors and beamforming weights. This is motivated by the fact that in real networks, only the beam indices and corresponding nominal beam directions are commonly available [32]. Furthermore, various practical aspects such as errors in the antenna element spacings as well as mutual coupling create anyway varying levels of uncertainties in the true antenna patterns – in particular in mmWave networks [33] where analog/RF beamforming dominates and the design and implementation of antenna elements and beamforming units is tedious. Additionally, the proposed channel parameter estimator that builds on the singular value decomposition (SVD) of the AoA-AoD BRSRP map is shown to have built-in robustness against the antenna sidelobes which is a clear additional benefit compared to the prior art in [15]. The proposed method is also compatible with 5G NR PRS signal structure and beam-sweaping procedures [34].
Secondly, we focus on advancing the robustness and tolerance of snapshot SLAM, compared to the prior art in [16, 28, 29], against practical measurement imperfections or outliers while also seek to improve the SLAM problem identifiability. These aspects have high importance in real mmWave deployment environments, where the amount of the available LoS/non-line-of-sight (NLoS) measurements can easily vary from measurement location to another [35]. Additionally, despite the advances in channel parameter estimation, measurement outliers are commonly occurring, e.g., due to the clutter, multi-bounce phenomenon and antenna sidelobes [5]. Specifically, we improve the snapshot SLAM identifiability via including an appropriate regularization term into the objective function that embeddes the prior information to the processing system. Additionally, a robust cost function is introduced to handle outliers originating, e.g., from false detections or clutter.
@@
  • We propose a new propagation path AoA/AoD estimation method utilizing standardized BRSRP measurements, alleviating the need for knowledge of the underlying antenna system beampatterns while offering controlled robustness against antenna sidelobes;
  • We propose a new snapshot SLAM algorithm that offers increased system identifiability and improved robustness against measurement outliers compared to prior-art;
  • We evaluate and benchmark the performance of the proposed methods in a realistic indoor environment through accurate ray-tracing as well as experimental measurement campaign, both carried out at the 60 GHz band;
  • We release the complete 60 GHz I/Q measurement data set as well as the corresponding processed channel parameter data set, together with supportive scripts for their utilization in any follow-up research;
@@ II-A Basic Assumptions and System Geometry
To this end, we consider transmitter (TX) and receiver (RX) entities equipped with uniform planar arrays (UPAs) with vertical times horizontal dimensions of the form ${N}^{v}_{\textrm{TX}}\times{N}^{h}_{\textrm{TX}}$ and ${N}^{v}_{\textrm{RX}}\times{N}_{\textrm{RX}}^{h}$, respectively. The antenna elements are separated by a half-wavelength distance $d=\lambda/2$, where $\lambda=\text{c}/f_{c}$ denotes the wavelength at carrier frequency $f_{c}$, and c is the speed of light. The total number of antenna elements in TX and RX are ${N}_{\textrm{TX}}={N}^{v}_{\textrm{TX}}\times{N}^{h}_{\textrm{TX}}$ and ${N}_{\textrm{RX}}={N}^{v}_{\textrm{RX}}\times{N}_{\textrm{RX}}^{h}$, respectively. The locations (3D coordinates) of the involved TX and RX entities are denoted by ${\mathbf{p}}_{\textrm{TX}}^{\textrm{3D}}=[{\mathbf{p}}_{\textrm{TX}}^{\top},\,z_{\textrm{TX}}]^{\top}$ and ${\mathbf{p}}_{\textrm{RX}}^{\textrm{3D}}=[{\mathbf{p}}_{\textrm{RX}}^{\top},\,z_{\textrm{RX}}]^{\top}$, respectively, where ${\mathbf{p}}_{\textrm{TX}}$ and ${\mathbf{p}}_{\textrm{RX}}$ are the corresponding 2D coordinates, all in global coordinate system. Additionally, the 3D location of an arbitrary single-bounce landmark is denoted by ${\mathbf{m}}_{n}^{\textrm{3D}}=[{\mathbf{m}}_{n}^{\top},\,z_{n}]^{\top}$ where ${\mathbf{m}}_{n}$ refers to the respective 2D coordinates while the subscript $n$ serves as a landmark or path index. Furthermore, we denote the 3D AoD as $\boldsymbol{\psi}^{\top}_{\text{TX},n}=[\theta_{n},\theta_{n}^{\text{EL}}]$, where $\theta_{n}$ and $\theta_{n}^{\text{EL}}$ are the azimuth and elevation AoDs, respectively, Similarly, the 3D AoA is denoted by $\boldsymbol{\psi}^{\top}_{\text{RX},n}=[\phi_{n},\phi_{n}^{\text{EL}}]$. All the involved angles are expressed in local coordinate systems, i.e., relative to the local orientations of the TX and RX entities.
@@ II-B Received Signal Model
where $\mathbf{w}_{\text{TX},i}\in\mathbb{C}^{\text{N}_{\textrm{TX}}}$ and $\mathbf{w}_{\text{RX},j}\in\mathbb{C}^{\text{N}_{\textrm{RX}}}$ are the TX and RX beamformers, $x_{k,m}\in\mathbb{C}$ with $|x_{k,m}|=1\,\forall k,m$ is the transmitted PRS sample, and $\mathbf{n}_{k,m}^{i,j}\in\mathbb{C}^{\text{N}_{RX}}$ denotes the antenna element wise additive white Gaussian noise (AWGN) at the RX. Furthermore, $\mathbf{H}_{k,m}\in\mathbb{C}^{\text{N}_{\textrm{RX}}\times\text{N}_{\textrm{TX}}}$ is the spatial multipath channel matrix defined as [35, eq. (4)]
$\begin{split}\mathbf{H}_{k,m}=\sum_{n=1}^{N}&\xi_{n}e^{-\mathrm{i}2\pi k\Delta f\tau_{n}}e^{\mathrm{i}2\pi mT_{\text{sym}}f_{\text{D},n}}\\ &\times\mathbf{a}_{\text{RX}}(\boldsymbol{\psi}_{\text{RX},n})\mathbf{a}_{\text{TX}}(\boldsymbol{\psi}_{\text{TX},n})^{\top},\end{split}$ (2)
where $\Delta f$ is the subcarrier spacing (SCS), $T_{\text{sym}}$ denotes the OFDM symbol duration, and $\xi_{n}$, $\tau_{n}$ and $f_{\text{D},n}$ are the complex path coefficient, the propagation delay, and the Doppler frequency for the $n^{\text{th}}$ propagation path, respectively. The delay $\tau_{n}$ maps to the corresponding propagation distance as $d_{n}=c\tau_{n}$. Additionally, $\mathbf{a}_{\text{TX}}(\boldsymbol{\psi}_{\text{TX},n})\!\in\!\mathbb{C}^{N_{\text{TX}}}$ and $\mathbf{a}_{\text{RX}}(\boldsymbol{\psi}_{\text{RX},n})\!\in\!\mathbb{C}^{N_{\text{RX}}}$ denote the TX and RX steering vectors, respectively. The exact way how the PRS sequences are mapped to the physical resources (OFDM symbols and the underlying subcarriers), in 5G NR context, is described in [34].
[…]
@@
$\begin{split}y_{k,m}^{i,j}=\Big{\{}\sum_{n=1}^{N}&\xi_{n}e^{-\mathrm{i}2\pi k\Delta f\tau_{n}}e^{\mathrm{i}2\pi mT_{\text{sym}}f_{\text{D},n}}x_{k,m}^{i,j}\\ &\times G_{\text{TX},i}(\boldsymbol{\psi}_{\text{TX},n})G_{\text{RX},j}(\boldsymbol{\psi}_{\text{RX},n})\Big{\}}+\tilde{n}_{k,m}^{i,j},\end{split}$ (3)

### v2

@@ Abstract
In this article, we address the timely topic of cellular bistatic simultaneous localization and mapping (SLAM) with specific focus on end-to-end processing solutions, from raw I/Q samples, via channel parameter estimation to user equipment (UE) and landmark location information in millimeter-wave (mmWave) networks, with minimal prior knowledge. Firstly, we propose a new multipath channel parameter estimation solution that operates directly with beam reference signal received power (BRSRP) measurements, alleviating the need to know the true antenna beampatterns or the underlying beamforming weights. Additionally, the method has built-in robustness against unavoidable antenna sidelobes. Secondly, we propose new snapshot SLAM algorithms that have increased robustness and identifiability compared to prior art, in practical built environments with complex clutter and multi-bounce propagation scenarios, and do not rely on any a priori motion model. The performance of the proposed methods is assessed at the 60 GHz mmWave band, via both realistic ray-tracing evaluations as well as true experimental measurements, in an indoor environment. A wide set of offered results demonstrate the improved performance, compared to the relevant prior art, in terms of the channel parameter estimation as well as the end-to-end SLAM performance. Finally, the article provides the measured 60 GHz data openly available for the research community, facilitating results reproducibility as well as further algorithm development.
@@ I Introduction
[…]
[lines=2]W  hile the primary purpose of mobile cellular networks is to provide efficient connectivity services, the ability to extract location information and situational awareness of the surrounding environment is also receiving increasing interest [1, 2, 3, 4]. The related notion of integrated sensing and communications (ISAC) refers to extending the situational awareness from ordinary user equipment (UE) positioning to the ability to sense also various passive objects in the environment, through cellular radio-based measurements and corresponding signal processing [5, 1]. Such knowledge of the UE locations and surrounding environment can be harnessed in numerous ways, for example in different extended reality (XR) use cases [6], vehicular applications [7], or industrial systems [8]. In general, the prospects for high-accuracy situational awareness are known to improve [9, 10] when the networks are expanding towards the millimeter-wave (mmWave) frequency bands. The current fifth generation New Radio (5G NR) specifications support already operating bands up to 71 GHz [11], while further extensions towards the sub-THz regime are expected in the 6G era [12].
@@
Bistatic cellular simultaneous localization and mapping (SLAM) is one of the prominent ISAC applications where the coordinates of both the UE as well as those of the environment scattering points – commonly referred to as the landmarks – are all jointly estimated, based on either uplink (UL) or downlink (DL) reference signals and known base-station (BS) locations [13, 14, 15]. This is illustrated conceptually in Fig. 1. Complete end-to-end solutions for SLAM comprise estimators for spatio-temporal channel parameters, often in the form of time of arrival (ToA), angle-of-arrival (AoA), and/or angle-of-departure (AoD) of the involved propagation paths, combined with the actual SLAM filters or snapshot estimators that process the channel parameter estimates into the corresponding estimates of the UE and landmark locations [16]. This is also the main scope of this article, in contrast to many other studies that focus either exclusively on channel parameter estimation in the context of wireless communication or on cellular SLAM, without considering the challenges of obtaining the channel parameters. We have a specific focus on implementation-feasible yet high-performing end-to-end processing solutions that can operate with minimal knowledge of the involved antenna system beampatterns and facilitate mmWave SLAM in practical complex built environments, particularly indoors, while operating on downlink positioning reference signal (PRS) standardized already for the existing 5G NR networks. Additionally, we emphasize research reproducibility and provide measured mmWave I/Q and channel data openly for the research community, while use the same measured data to evaluate and benchmark the proposed methods against the relevant prior art.
@@ I-A Prior Art
The available literature on channel parameter estimation is generally-speaking wide, however, a vast majority of works is carried out under the assumptions of known antenna steering vectors, beamforming weights, and thereon beampatterns. Such methods are described, e.g., in [17, 18, 19, 20, 21, 22, 23, 24, 25]. To this end, [17] and [18] present Bayesian channel estimation algorithms based on different variants of Kalman filters, while [19, 21, 22] harness the mmWave channel sparsity through compressed sensing (CS) methods. The work in [20], in turn, proposes an AoA estimation method using virtual subarrays which, however, requires a special beamformer or antenna pattern design, similar to [23]. Furthermore, [24, 25] propose joint AoD and AoA estimation methods with specific frequency-dependent codebooks in true-time-delay array context. Opposed to these previous works, an angle estimation algorithm building on the beam reference signal received power (BRSRP) measurements is described in the recent work in [15]. The method builds on thresholding and successive cancellation principle, operating on AoA-AoD power map, however, is lacking, e.g., explicit treatment of antenna sidelobes whose impact can be substantial in mmWave systems. This is particularly so in SLAM context where the multipaths and their dynamic range are one of the key factors. For clarity, we also note that works like [26] exist that deploy machine-learning to the channel parameters estimation – however, such works are still commonly at their early phases.
@@
In the context of the available SLAM methods, we first note that both snapshot approaches [16, 27, 28, 29] as well as sequential filtering based solutions [30, 13, 14, 15, 31] exist. Both research directions are relevant and the preferred choice depends on the overall system and application scenario. Snapshot SLAM is fundamentally important as it serves as a baseline for what can be done with radio signals alone, without any movement models, while a snapshot method can also be used as input to filtering [27]. On the other hand, filtering based SLAM methods process the observations sequentially over time and are expected to remove false detections and improve the accuracy [14]. Both SLAM methods also have their limitations. A major drawback of snapshot SLAM is that it is not always applicable, since the measurements may not be sufficient to solve the SLAM problem [16]. For example, the UE cannot be localized with the line-of-sight (LoS) alone if the clocks of the UE and BS are not synchronized [27] and one must resort to filtering based techniques to solve the problem over time. On the other hand, filtering methods always require a snapshot algorithm for initialization when prior information is not available. Another disadvantage of filtering methods is that they require solving a complex data association problem which increases computational overhead of the algorithms [14, 31]. It is to be noted that low complexity alternatives exist [30, 13], however, at the expense of reduced accuracy. Finally, different from mono-static sensing where transmitter (TX) and receiver (RX) entities are directly mutually synchronized [5],[32, 33], an important practical aspect and challenge in bi-static SLAM is the ability to estimate and track the time-varying clock-bias between the UE and the network elements. Examples of existing works where clock parameters are incorporated as part of the overall state estimation or tracking problem include, e.g., [27, 14].
@@ I-B Novelty and Contributions
Compared to the available methods and literature, this article provides the following contributions. Firstly, as opposed to the vast majority of existing literature in [17, 18, 19, 20, 21, 22, 23, 24, 25], we focus on enabling accurate multipath AoA/AoD estimation with BRSRP measurements only, without knowledge of the complex antenna patterns or the underlying steering vectors and beamforming weights. This is motivated by the fact that in real networks, only the beam indices and corresponding nominal beam directions are commonly available [34]. Furthermore, various practical aspects such as errors in the antenna element spacings as well as mutual coupling create anyway varying levels of uncertainties in the true antenna patterns – in particular in mmWave networks [35] where analog/RF beamforming dominates and the design and implementation of antenna elements and beamforming units are tedious. Additionally, the proposed channel parameter estimator that builds on the singular value decomposition (SVD) of the AoA-AoD BRSRP map is shown to have built-in robustness against the antenna sidelobes which is a clear additional benefit compared to the prior art in [15]. The proposed method is also compatible with 5G NR PRS signal structure and beam-sweaping procedures defined in [36].
Secondly, we focus on advancing the SLAM in three directions: first, through increased robustness and tolerance of snapshot SLAM, compared to the prior art in [16, 28, 29], against practical measurement imperfections or outliers; second, by improving the SLAM problem identifiability; and third, to perform SLAM without knowledge of the user motion model. These aspects have high importance in real mmWave deployment environments, where the amount of the available LoS/non-line-of-sight (NLoS) measurements can easily vary from measurement location to another [37]. Additionally, despite the advances in channel parameter estimation, measurement outliers are commonly occurring, e.g., due to the clutter, multi-bounce phenomenon and antenna sidelobes [5]. Specifically, we improve the snapshot SLAM identifiability via including an appropriate regularization term into the objective function that embeddes the prior information to the processing system. Additionally, a robust cost function is introduced to handle outliers originating, e.g., from false detections or clutter.
@@
  • End-to-end robust snapshot SLAM approach: We develop and demonstrate an approach for end-to-end snapshot SLAM, under minimal knowledge of the array parameters and the user motion model. The end-to-end approach comprises a novel channel parameter estimation method and a novel snapshot SLAM method.
  • Robust channel parameter estimation: We propose a new propagation path AoA/AoD estimation method utilizing standardized BRSRP measurements, alleviating the need for knowledge of the underlying antenna system beampatterns while offering controlled robustness against antenna sidelobes;
  • Robust snapshot SLAM: We propose a new snapshot SLAM algorithm that offers increased system identifiability and improved robustness against measurement outliers compared to prior-art;
  • Realistic performance evaluation: We evaluate and benchmark the performance of the proposed methods in a realistic indoor environment through accurate ray-tracing as well as experimental measurement campaign, both carried out at the 60 GHz band;
  • Open-source data-set: We release the complete 60 GHz I/Q measurement data set as well as the corresponding processed channel parameter data set, together with supportive scripts for their utilization in any follow-up research;
@@ II-A Basic Assumptions and System Geometry
To this end, we consider TX and RX entities equipped with uniform planar arrays (UPAs) with vertical times horizontal dimensions of the form ${N}^{v}_{\textrm{TX}}\times{N}^{h}_{\textrm{TX}}$ and ${N}^{v}_{\textrm{RX}}\times{N}_{\textrm{RX}}^{h}$, respectively. The antenna elements are separated by a half-wavelength distance $d=\lambda/2$, where $\lambda=c/f_{\text{c}}$ denotes the wavelength at carrier frequency $f_{\text{c}}$, and $c$ is the speed of light. The total number of antenna elements in TX and RX are ${N}_{\textrm{TX}}={N}^{v}_{\textrm{TX}}\times{N}^{h}_{\textrm{TX}}$ and ${N}_{\textrm{RX}}={N}^{v}_{\textrm{RX}}\times{N}_{\textrm{RX}}^{h}$, respectively. The locations (3D coordinates) of the involved TX and RX entities are denoted by ${\mathbf{p}}_{\textrm{TX}}^{\textrm{3D}}=[{\mathbf{p}}_{\textrm{TX}}^{\top},\,z_{\textrm{TX}}]^{\top}$ and ${\mathbf{p}}_{\textrm{RX}}^{\textrm{3D}}=[{\mathbf{p}}_{\textrm{RX}}^{\top},\,z_{\textrm{RX}}]^{\top}$, respectively, where ${\mathbf{p}}_{\textrm{TX}}$ and ${\mathbf{p}}_{\textrm{RX}}$ are the corresponding horizontal 2D coordinates, while $z_{\textrm{TX}}$ and $z_{\textrm{RX}}$ denote the respective vertical coordinates – all in global coordinate system. Additionally, the 3D location of an arbitrary single-bounce landmark is denoted by ${\mathbf{m}}_{n}^{\textrm{3D}}=[{\mathbf{m}}_{n}^{\top},\,z_{n}]^{\top}$, where ${\mathbf{m}}_{n}$ refers to the respective horizontal 2D coordinates, $z_{n}$ is a vertical coordinate, while the subscript $n$ serves as a landmark or path index. Furthermore, we denote the 3D AoD as $\boldsymbol{\psi}^{\top}_{\text{TX},n}=[\phi_{n},\phi_{n}^{\text{EL}}]$, where $\phi_{n}$ and $\phi_{n}^{\text{EL}}$ are the azimuth and elevation AoDs, respectively, Similarly, the 3D AoA is denoted by $\boldsymbol{\psi}^{\top}_{\text{RX},n}=[\theta_{n},\theta_{n}^{\text{EL}}]$. All the involved angles are expressed in local coordinate systems, i.e., relative to the local orientations of the TX and RX entities.
@@ II-B Received Signal Model
where $\mathbf{w}_{\text{TX},i}\in\mathbb{C}^{\text{N}_{\textrm{TX}}}$ and $\mathbf{w}_{\text{RX},j}\in\mathbb{C}^{\text{N}_{\textrm{RX}}}$ are the TX and RX beamformers, $x_{k,m}\in\mathbb{C}$ with $|x_{k,m}|=1\,\forall k,m$ is the transmitted PRS sample, and $\mathbf{n}_{k,m}^{i,j}\in\mathbb{C}^{\text{N}_{RX}}$ denotes the antenna element wise additive white Gaussian noise (AWGN) at the RX. Furthermore, $\mathbf{H}_{k,m}\in\mathbb{C}^{\text{N}_{\textrm{RX}}\times\text{N}_{\textrm{TX}}}$ is the effective spatial multipath channel matrix defined as [37, eq. (4)]
$\begin{split}\mathbf{H}_{k,m}=\sum_{n=1}^{N}&\xi_{n}e^{-\mathrm{i}2\pi k\Delta f{\tau_{\mathrm{f},n}}}e^{\mathrm{i}2\pi mT_{\text{sym}}f_{\text{D},n}}\\ &\times\mathbf{a}_{\text{RX}}(\boldsymbol{\psi}_{\text{RX},n})\mathbf{a}_{\text{TX}}(\boldsymbol{\psi}_{\text{TX},n})^{\top},\end{split}$ (2)
where $\Delta f$ is the subcarrier spacing (SCS), $T_{\text{sym}}$ denotes the OFDM symbol duration, and $\xi_{n}$, $\tau_{\mathrm{f},n}$ and $f_{\text{D},n}$ are the complex path coefficient, the propagation delay with respect to the beginning of the received OFDM symbol, and the Doppler frequency for the $n^{\text{th}}$ propagation path, respectively. The true physical propagation delay for path $n$ is denoted by $\tau_{n}=\tau_{\text{c}}+\tau_{\mathrm{f},n}$, with $d_{n}=c\tau_{n}$ denoting the corresponding physical distance, where $\tau_{\text{c}}$ refers to the delay between the received OFDM symbol and transmit time. Considering now a receiver with an unknown clock bias $b_{\textrm{UE}}$, the corresponding biased propagation delay $\tau^{b}_{n}$ can be expressed as
$\tau^{b}_{n}=\tau_{n}-b_{\textrm{UE}}=\tau_{\text{c}}+\tau_{\mathrm{f},n}-b_{\textrm{UE}}.$ (3)
Estimation of $\tau_{\text{c}}$, subject to unknown bias $b_{\textrm{UE}}$, can be considered part of regular OFDM symbol synchronization, for example, through time-based correlation, as discussed further in Section III-C. Additionally, $\mathbf{a}_{\text{TX}}(\boldsymbol{\psi}_{\text{TX},n})\!\in\!\mathbb{C}^{N_{\text{TX}}}$ and $\mathbf{a}_{\text{RX}}(\boldsymbol{\psi}_{\text{RX},n})\!\in\!\mathbb{C}^{N_{\text{RX}}}$ denote the TX and RX steering vectors, respectively. The exact way how the PRS sequences are mapped to the physical resources (OFDM symbols and the underlying subcarriers), in 5G NR context, is described in [36].
@@
$\begin{split}y_{k,m}^{i,j}=\Big{\{}\sum_{n=1}^{N}&\xi_{n}e^{-\mathrm{i}2\pi k\Delta f{\tau_{\mathrm{f},n}}}e^{\mathrm{i}2\pi mT_{\text{sym}}f_{\text{D},n}}x_{k,m}^{i,j}\\ &\times G_{\text{TX},i}(\boldsymbol{\psi}_{\text{TX},n})G_{\text{RX},j}(\boldsymbol{\psi}_{\text{RX},n})\Big{\}}+\tilde{n}_{k,m}^{i,j},\end{split}$ (4)